# BARD: Budget-Aware Reasoning Distillation

### v1

@@ 1 Introduction
Large Reasoning Models (LRMs) have demonstrated remarkable capabilities in complex reasoning tasks that require multi-step logic and structured problem-solving [1, 2, 3]. Long Chain-of-Thought (CoT) reasoning [4] significantly improves both the performance and interpretability of LLMs by encouraging them to articulate intermediate reasoning steps. However, state-of-the-art reasoning models typically possess enormous parameter scales, making their deployment and inference prohibitively expensive in long CoT reasoning modes. Such costs severely hinder their use in real-world applications, particularly in latency-sensitive or resource-constrained environments.
To address this challenge, knowledge distillation, and specifically CoT distillation [5, 6] , has emerged as a popular technique for transferring the superior reasoning abilities of large teacher models to smaller and more efficient student models. However, the length of the generated reasoning chain remains uncontrolled and often just inherits the teacher’s (frequently verbose) reasoning style, leading to unstable computational costs and inefficient resource utilization. In addition, the distribution of generated reasoning lengths exhibits a natural long-tail pattern, which poses challenges in achieving fine-grained control over reasoning length across a diverse range [7, 8].
@@ 2 Related Work
Recently, Large Reasoning Models (LRMs) have demonstrated impressive reasoning capabilities at large scales, while smaller models often require knowledge distillation to achieve comparable performance [1]. However, while distillation effectively enhances reasoning capability, it offers limited control over the reasoning length, often leading to uncontrollable or inefficient reasoning behaviors.
To address this limitation, the most straightforward approach to budget control is budget forcing  [9], which simply truncates a model’s thinking process to a desired length. While easy to implement, these naive methods are non-selective and risk severing critical logical steps, often leading to a catastrophic collapse in reasoning performance.A more sophisticated line of work uses supervised fine-tuning (SFT) on specially curated datasets. These methods aim to teach the model a specific reasoning style or length preference. C3oT [10] enables users to switch between short and long reasoning modes, providing coarse-grained control over CoT length. [11] constructs a fine-tuning dataset composed of the shortest correct reasoning paths to encourage concise reasoning. However, such methods offer only discrete control (short vs. long), limiting user flexibility. TokenSkip [12] allows users to specify the CoT length as a ratio of the original reasoning length, skipping unimportant tokens to generate shorter CoTs, though this approach often produces fragmented and less interpretable reasoning.
In contrast, other works directly target reasoning models, leveraging reinforcement learning to regulate reasoning length. Methods such as L1 [13], O1-Pruner [14], and [15] introduce length-dependent rewards to constrain reasoning generation. While effective to some extent, these methods face exploration inefficiency due to the vast output space and often fail to learn a stable mapping between user-specified length control and internal reasoning dynamics.
@@ 4.1 Experimental Setup
We use Qwen38B-Base as our student model throughout all experiments. The teacher model, DeepSeek-R1[1], is employed to generate the initial long-form Chain-of-Thought (CoT) rationales. Our training methodology relies on a vast and heterogeneous mixture of long Chain-of-Thought (CoT) data to ensure broad generalization. The dataset for the Supervised Fine-Tuning (SFT) phase is a comprehensive blend of multi-domain sources, sampled from a combination of publicly available open-source datasets [16, 17, 18] and in-house data sources, can be categorized as:
@@
First, we determine the length of the original, full-length CoT, denoted as $L_{full}$. We then randomly sample three distinct budget values, $\{b_{1},b_{2},b_{3}\}$, from the range $0,L_{full}]$ to serve as predefined target lengths. For each sampled budget $b_{i}$, we employ a powerful teacher model, Qwen3-32B [19], to compress the original CoT to be within this budget.
In the subsequent Reinforcement Learning (RL) phase, we shift to a more targeted data mixture to refine specific abilities. We construct the RL training set by combining open-ended QA data [17, 18] and mathematical reasoning data [20] in a 1:1 ratio. To simulate diverse user requirements and enhance the model’s adaptability, the user-specified budget $b$ for each training instance is dynamically sampled from a predefined range during training.
@@
We first perform Supervised Fine-Tuning (SFT) on a mixed 1300K dataset (390K budget-aware, 350K standard CoT) for 4 epochs with a learning rate of 1.0e-5. For the subsequent Reinforcement Learning (RL) phase, we use a mixed 10K dataset and employ the Group-wise Ranking Proximal Optimization (GRPO) [21] algorithm. Our implementation is built upon ROLL [22]. We sample 16 responses per group with a batch size of 64. The GRPO difficulty thresholds are set to 0.1 (low) and 0.95 (high) to filter comparison pairs, and the learning rate is 1.0e-6.
@@ 4.2 Evaluation Metrics
We evaluate all models on the AIME 2024, AIME 2025, and GPQA [23] benchmarks. Budget-aware models are tested across a wide range of budgets: $b\in\{500,1000,1500,3000,5000,8000\}$ tokens. We use two primary metrics:
@@
[24] To provide a holistic score for model comparison, we introduce a Unified Performance Score (UPS). This metric combines both accuracy and budget fidelity into a single value, reflecting the trade-offs made by different models. It is defined as a weighted sum:
@@ 4.4 RQ2: Ablation Studies on BARD Components
We first evaluate the importance of our initial, budget-aware SFT phase. To do this, we use an ablated variant, RL-on-SFT-Full, where RL is applied directly to a checkpoint that was only fine-tuned on long CoTs without any budget information (the SFT-Full model). This setup is conceptually similar to some prior distillation approaches [13]. As shown in Figure 3, the resulting model was completely unresponsive to the thinking budget command and exhibited no length control capabilities. This confirms our hypothesis that our budget aware SFT phase is a critical bootstrapping step. It is indispensable for instilling an initial understanding of the budget command’s meaning, providing a foundational policy that makes the subsequent RL exploration tractable. Without it, the model faces an intractably large exploration space, leading to policy collapse. Table 1 presents a comprehensive comparison of UPS scores. The results demonstrate BARD’s ability to reason more efficiently, leading to a superior balance of performance and cost-effectiveness.
Value of the contrastive learning.
@@ 4.5 RQ3: Analysis of Reasoning Behavior Change
Figure 7: Analysis of BARD’s Adaptive Reasoning Process under Different Budgets. The figure visualizes the distribution of reasoning step types generated by BARD under a low-budget constraint (500 tokens, left) versus a high-budget constraint (8000 tokens, right). Each color represents a different category [25] of reasoning step. The change in the relative proportions of these steps illustrates how the model strategically allocates its cognitive resources depending on the available budget.

### v2

@@ 1 Introduction
Large Reasoning Models (LRMs) have demonstrated remarkable capabilities in complex reasoning tasks that require multi-step logic and structured problem-solving [guo2025deepseek, jaech2024openai, yang2025qwen3]. Long Chain-of-Thought (CoT) reasoning [wei2022chain] significantly improves both the performance and interpretability of LLMs by encouraging them to articulate intermediate reasoning steps. However, state-of-the-art reasoning models typically possess enormous parameter scales, making their deployment and inference prohibitively expensive in long CoT reasoning modes. Such costs severely hinder their use in real-world applications, particularly in latency-sensitive or resource-constrained environments.
To address this challenge, knowledge distillation, and specifically CoT distillation [wang2025efficient, niu-etal-2025-cotd] , has emerged as a popular technique for transferring the superior reasoning abilities of large teacher models to smaller and more efficient student models. However, the length of the generated reasoning chain remains uncontrolled and often just inherits the teacher’s (frequently verbose) reasoning style, leading to unstable computational costs and inefficient resource utilization. In addition, the distribution of generated reasoning lengths exhibits a natural long-tail pattern, which poses challenges in achieving fine-grained control over reasoning length across a diverse range [kandpal2023large, zhang2023deep].
@@ 2 Related Work
Recently, Large Reasoning Models (LRMs) have demonstrated impressive reasoning capabilities at large scales, while smaller models often require knowledge distillation to achieve comparable performance [guo2025deepseek]. However, while distillation effectively enhances reasoning capability, it offers limited control over the reasoning length, often leading to uncontrollable or inefficient reasoning behaviors.
To address this limitation, the most straightforward approach to budget control is budget forcing  [muennighoff2025s1], which simply truncates a model’s thinking process to a desired length. While easy to implement, these naive methods are non-selective and risk severing critical logical steps, often leading to a catastrophic collapse in reasoning performance.A more sophisticated line of work uses supervised fine-tuning (SFT) on specially curated datasets. These methods aim to teach the model a specific reasoning style or length preference. C3oT [kang2025c3ot] enables users to switch between short and long reasoning modes, providing coarse-grained control over CoT length. [munkhbat2025self] constructs a fine-tuning dataset composed of the shortest correct reasoning paths to encourage concise reasoning. However, such methods offer only discrete control (short vs. long), limiting user flexibility. TokenSkip [xia2025tokenskip] allows users to specify the CoT length as a ratio of the original reasoning length, skipping unimportant tokens to generate shorter CoTs, though this approach often produces fragmented and less interpretable reasoning.
In contrast, other works directly target reasoning models, leveraging reinforcement learning to regulate reasoning length. Methods such as L1 [aggarwal2025l1], O1-Pruner [luo2025o1], and [arora2025training] introduce length-dependent rewards to constrain reasoning generation. While effective to some extent, these methods face exploration inefficiency due to the vast output space and often fail to learn a stable mapping between user-specified length control and internal reasoning dynamics.
@@ 4.1 Experimental Setup
We use Qwen38B-Base as our student model throughout all experiments. The teacher model, DeepSeek-R1[guo2025deepseek], is employed to generate the initial long-form Chain-of-Thought (CoT) rationales. Our training methodology relies on a vast and heterogeneous mixture of long Chain-of-Thought (CoT) data to ensure broad generalization. The dataset for the Supervised Fine-Tuning (SFT) phase is a comprehensive blend of multi-domain sources, sampled from a combination of publicly available open-source datasets [numina_math_datasets, park2025instruct, lambert2024tulu] and in-house data sources, can be categorized as:
@@
First, we determine the length of the original, full-length CoT, denoted as $L_{full}$. We then randomly sample three distinct budget values, $\{b_{1},b_{2},b_{3}\}$, from the range $0,L_{full}]$ to serve as predefined target lengths. For each sampled budget $b_{i}$, we employ a powerful teacher model, Qwen3-32B [qwen3technicalreport], to compress the original CoT to be within this budget.
In the subsequent Reinforcement Learning (RL) phase, we shift to a more targeted data mixture to refine specific abilities. We construct the RL training set by combining open-ended QA data [park2025instruct, lambert2024tulu] and mathematical reasoning data [luo2025deepscaler] in a 1:1 ratio. To simulate diverse user requirements and enhance the model’s adaptability, the user-specified budget $b$ for each training instance is dynamically sampled from a predefined range during training.
@@
We first perform Supervised Fine-Tuning (SFT) on a mixed 1300K dataset (390K budget-aware, 350K standard CoT) for 4 epochs with a learning rate of 1.0e-5. For the subsequent Reinforcement Learning (RL) phase, we use a mixed 10K dataset and employ the Group-wise Ranking Proximal Optimization (GRPO) [shao2024deepseekmath] algorithm. Our implementation is built upon ROLL [wang2025reinforcement]. We sample 16 responses per group with a batch size of 64. The GRPO difficulty thresholds are set to 0.1 (low) and 0.95 (high) to filter comparison pairs, and the learning rate is 1.0e-6.
@@ 4.2 Evaluation Metrics
We evaluate all models on the AIME 2024, AIME 2025, and GPQA [rein2024gpqa] benchmarks. Budget-aware models are tested across a wide range of budgets: $b\in\{500,1000,1500,3000,5000,8000\}$ tokens. We use two primary metrics:
@@
[he2025thinkdial] To provide a holistic score for model comparison, we introduce a Unified Performance Score (UPS). This metric combines both accuracy and budget fidelity into a single value, reflecting the trade-offs made by different models. It is defined as a weighted sum:
@@ 4.4 RQ2: Ablation Studies on BARD Components
We first evaluate the importance of our initial, budget-aware SFT phase. To do this, we use an ablated variant, RL-on-SFT-Full, where RL is applied directly to a checkpoint that was only fine-tuned on long CoTs without any budget information (the SFT-Full model). This setup is conceptually similar to some prior distillation approaches [aggarwal2025l1]. As shown in Figure 3, the resulting model was completely unresponsive to the thinking budget command and exhibited no length control capabilities. This confirms our hypothesis that our budget aware SFT phase is a critical bootstrapping step. It is indispensable for instilling an initial understanding of the budget command’s meaning, providing a foundational policy that makes the subsequent RL exploration tractable. Without it, the model faces an intractably large exploration space, leading to policy collapse. Table 1 presents a comprehensive comparison of UPS scores. The results demonstrate BARD’s ability to reason more efficiently, leading to a superior balance of performance and cost-effectiveness.
@@ 4.5 RQ3: Analysis of Reasoning Behavior Change
Figure 7: Analysis of BARD’s Adaptive Reasoning Process under Different Budgets. The figure visualizes the distribution of reasoning step types generated by BARD under a low-budget constraint (500 tokens, right) versus a high-budget constraint (8000 tokens, left). Each color represents a different category [hou2025thinkprune] of reasoning step. The change in the relative proportions of these steps illustrates how the model strategically allocates its cognitive resources depending on the available budget.
@@ 5 Conclusion
Appendix A Prompt for Thought Process Compression
This appendix details the prompt used to instruct the LLM to compress its thought process. The placeholders {query}, {think}, {answer}, and {think_len} were dynamically filled during the process.
## Task Description
You are an expert in compressing "Thought Processes." Please compress the
provided "Thought Process" according to the following requirements:
1. Refer to the input information below (the related Question, Thought
   Process, and Answer). You must analyze the relationship between the
   Question and the Answer, and compress the Thought Process to a
   specified length. It is crucial that you do not alter the original style
   or meaning of the thought process. The compressed thought process must
   serve as a logical bridge between the Question and the Answer, ensuring
   coherence.
2. While compressing the Thought Process to the specified token limit,
   avoid excessive compression. Strive to retain the most critical content
   of the thought process.
3. The first sentence of the original Thought Process must remain
   unchanged.
## Input Information
### Question
{query}
### Thought Process
{think}
### Answer
{answer}
## Output Requirements
1. Compressed Thought Process Length: {think_len} tokens
2. Output format for the compressed thought process:
   Compressed Thought Process:(Ensure no other content follows)